# DyBBT: Dynamic Balance via Bandit inspired Targeting for Dialog Policy with Cognitive Dual-Systems

### v1

@@ D.4 Hyperparameter Sensitivity Analysis
D.5 Model Scaling Analysis
@@
To systematically evaluate the impact of model scale on DyBBT’s performance and efficiency, we conduct a comprehensive scaling analysis using three prominent open weight model families: Llama-3.2 Instruct(1B–8B), Qwen2.5 Instruct(0.5B–7B), and Qwen3 (0.6B–8B) on the MultiWOZ 2.1 benchmark. Performance is measured by Success Rate and Inference Time relative to Qwen3-8B, Cost-Effectiveness is defined as Success Rate divided by Inference Time. Results are summarized in Table 6.
@@
These performance gains come with increased computational cost. Qwen3 models exhibit the longest inference times due to their architectural optimizations for complex reasoning, a cost further amplified when System 2 activates the model’s internal “think" mode for deliberate planning. Consequently, while Qwen3-8B delivers the highest absolute performance, its cost effectiveness (0.851) is lower than that of smaller models. Among the larger models, Qwen2.5-7B offers a favorable balance, achieving 97.6% of the performance of Qwen3-8B at 86% of the inference cost.
@@
D.6 Comparison with Qwen3’s Native Switching
Table 7: Comparison between DyBBT’s meta-controller and Qwen3’s native switching mechanism. Inference time is normalized to DyBBT’s default mode (S1 no think / S2 think = 1.0x).
Configuration Success Rate $\uparrow$ Inference Time $\downarrow$ Cost-Effectiveness $\uparrow$ S1 no think / S2 no think $79.6\pm 0.015$ 0.6x 132.7 S1 think / S2 think $\mathbf{86.5\pm 0.010}$ 3.2x 27.0 DyBBT (S1 no think / S2 think) $85.1\pm 0.011$ 1.0x 85.1
@@ D.6 Comparison with Qwen3’s Native Switching
The no-think baseline performs poorly, underscoring the necessity of deliberate reasoning in complex dialog states. DyBBT strikes a balance between these extremes by invoking costly reasoning only when cognitively justified, either due to under exploration or low confidence, leading to near optimal performance with moderate and targeted computational overhead. This leads to less efficient allocation of computational resources, as also reflected in human evaluation (Section 4.4).
D.7 Case Study
@@
D.7.1 Case 1: Successful Intervention due to High Epistemic Uncertainty
@@
Belief State:
restaurant {
semi {
food: "Chinese" # (USER_CONFIRMED)
pricerange: "cheap" # (USER_CONFIRMED)
area: "" # (USER_MENTIONED but NOT_CONFIRMED)
name: "" # (NOT_MENTIONED - High Uncertainty)
book { people: "", day: "", time: "" }
taxi {
semi {
destination: "", departure: "", leaveAt: "", arriveBy: ""
@@
  • Slot Dependency ($\rho_{t}$): 0.6 (Medium, ‘area‘ and ‘name‘ often co-occur in restaurant domain)
Meta-Controller Decision: The visitation count for this cognitive state region was low ($n_{t}(\mathbf{c}_{t})=12<\tau\sqrt{\log T}\approx 25$), triggering System 2 via the exploration condition. System 1’s confidence was medium ($p_{t}^{S1}=0.75>\kappa$).
System 2 Intervention: System 2 performed multi-path reasoning. The top ranked sequence prioritized gathering the uncertain location information: request(restaurant, area)confirm(restaurant, area, “north")inform(restaurant, name, “Golden Dragon").
Outcome: This strategy efficiently disambiguated the user’s intent. The dialog was successfully completed 6 turns later. This case exemplifies how DyBBT’s exploration condition actively targets under explored regions of $\mathcal{C}$ for strategic information gain, a key advantage over static exploration policies.
D.7.2 Case 2: Successful Intervention due to Low Aleatoric Confidence
@@
Belief State:
train {
semi {
destination: "cambridge", # (CONFIRMED)
departure: "london", # (CONFIRMED)
day: "today" # (CONFIRMED)
book { people: "2" } # (CONFIRMED)
taxi {
semi {
destination: "london stansted airport", # (CONFIRMED - but contradicts train?)
departure: "cambridge" # (CONFIRMED)
@@
Meta-Controller Decision: The cognitive state was well explored ($n_{t}(\mathbf{c}_{t})=505$). However, System 1’s confidence was low ($p_{t}^{S1}=0.55<\kappa=0.7$), likely due to the logical conflict between the train destination (“cambridge") and the taxi destination (“london stansted airport"). This triggered System 2 via the confidence condition.
System 2 Intervention: System 2 identified the inconsistency. Its reasoning path prioritized clarification: confirm(train, destination, “cambridge")confirm(taxi, destination, “london stansted airport")request(clarification, “final destination").
Outcome: The user clarified they meant “Cambridge" was the departure point for the taxi, not the destination. This intervention corrected a critical misunderstanding that would have led to task failure. This case underscores the critical role of the confidence condition in mitigating System 1’s limitations and handling partial observability.
D.7.3 Case 3: Failure due to Cognitive State Misrepresentation
@@
Belief State:
hotel {
semi {
name: "hilton", # (CONFIRMED)
area: "centre", # (CONFIRMED)
parking: "yes", # (CONFIRMED)
pricerange: "expensive" # (CONFIRMED)
book { people: "2", day: "today", stay: "2 nights" } # (BOOKED)
attraction {
semi {
type: "museum", # (USER_MENTIONED)
name: "" # (NOT_MENTIONED)
area: "centre" # (INFERRED from hotel)
@@
Meta-Controller Decision: The state had medium visitation ($n_{t}(\mathbf{c}_{t})=162$) and System 1 was highly confident ($p_{t}^{S1}=0.92$) in its action to request(attraction, name). The meta-controller did not trigger System 2.
Analysis of Failure: While the cognitive state suggested a routine information gathering context, it failed to capture the user had just finished a complex booking and was likely expecting a concise recommendation, not another request. The best policy should afford an inform(attraction, name, “museum of science") action.
Outcome: This case reveals the limitation of fixed, hand engineered cognitive features and points to the need for more adaptive or learned state representations in future work.
D.7.4 Summary and Limitations
@@
Appendix E LLM Use Statement
We utilized DeepSeek V3.1 for translation assistance and grammatical refinement of certain textual passages, and employed Qwen3-Code to aid in debugging and optimizing portions of the experimental code. These LLMs served solely as support tools for improving linguistic clarity and technical implementation. They played no role in the conceptualization of the research, the formulation of methodologies, the analysis of results, or the derivation of scientific conclusions. Consequently, their use does not qualify them as contributors under the authorship criteria. The authors assume full responsibility for all aspects of the work, including the accuracy and integrity of all generated and modified content, and affirm that appropriate measures have been taken to prevent plagiarism and other forms of scientific misconduct.

### v2

@@ E.5 Hyperparameter Sensitivity Analysis
E.6 Model Scaling Analysis
@@
To systematically evaluate the impact of model scale on DyBBT’s performance and efficiency, we conduct a comprehensive scaling analysis using three prominent open weight model families: Llama-3.2 Instruct(1B-8B), Qwen2.5 Instruct(0.5B-7B), and Qwen3 (0.6B-8B) on the MultiWOZ 2.1 benchmark. Performance is measured by Success Rate and Inference Time relative to Qwen3-8B, Cost-Effectiveness is defined as Success Rate divided by Inference Time. Results are summarized in Table 8.
@@
These performance gains come with increased computational cost. Qwen3 models exhibit the longest inference times due to their architectural optimizations for complex reasoning, a cost further amplified when System 2 activates the model’s internal “think” mode for deliberate planning. Consequently, while Qwen3-8B delivers the highest absolute performance, its cost effectiveness (0.851) is lower than that of smaller models. Among the larger models, Qwen2.5-7B offers a favorable balance, achieving 97.6% of the performance of Qwen3-8B at 86% of the inference cost.
@@
E.7 Cost Effectiveness Analysis of Different System Configurations
To provide practitioners with a clear cost performance trade off analysis, we compare DyBBT-8B, DyBBT-8B/GPT-4.0, and LLM_DP (pure GPT-4.0) on the MultiWOZ dataset. Since GPT-4.0 is only available via commercial APIs, we adopt two alternative evaluation approaches: measuring end-to-end inference time under the same hardware environment, and calculating economic cost based on actual token usage.
All local models run on an NVIDIA 5090 GPU, while the API model (GPT-4.0) is accessed via the official interface. The end-to-end Inference Time including model forward propagation or API call latency, averaged seconds per dialog. Normalized Inference Time is benchmarked against DyBBT-8B’s inference time. API Cost is based on GPT-4.0’s official pricing (input: \$0.03 per 1k tokens; output: \$0.06 per 1k tokens).
[…]
Table 9: Cost effectiveness analysis of different system configurations
Table 9 presents the comprehensive cost effectiveness comparison. Compared to DyBBT-8B, DyBBT-8B/GPT-4.0 achieves only a 1.2% improvement in success rate, but incurs a 2.3× increase in inference time and a cost of \$0.16 per dialog. This indicates that marginal performance gains are accompanied by substantial computational overhead and economic costs. LLM_DP (GPT-4.0), which relies solely on well designed prompts to enable LLMs to generate system actions, not only achieves an extremely low success rate but also has the longest inference time and highest API cost, highlighting the advantage of the DyBBT framework in balancing performance and cost. The System 2 invocation ratio of DyBBT-8B/GPT-4.0 is only 14.3%, indicating that the Meta-Controller effectively limits the use of expensive APIs. However, API call latency still dominates the total inference time.
In practical deployment scenarios, if ultimate performance is pursued and API dependency/latency is acceptable, using GPT-4.0 or more advanced closed source models for System 2 is an option. This requires balancing the 1.2% performance gain against the 2.3× inference time and additional costs. Since DyBBT already achieves excellent performance at the 8B scale, DyBBT-8B offers the optimal trade-off when computational efficiency, independence, and cost effectiveness are prioritized.
E.8 Comparison with Qwen3’s Native Switching
Configuration Success Rate $\uparrow$ Normalized Time $\downarrow$ Cost Effectiveness $\uparrow$ S1 no think / S2 no think $79.6\pm 0.015$ 0.6x 132.7 S1 think / S2 think $\mathbf{86.5\pm 0.010}$ 3.2x 27.0 DyBBT (S1 no think / S2 think) $85.1\pm 0.011$ 1.0x 85.1
Table 10: Comparison between DyBBT’s meta-controller and Qwen3’s native switching mechanism. Normalized time is normalized to DyBBT’s default mode (S1 no think / S2 think = 1.0x).
@@ E.8 Comparison with Qwen3’s Native Switching
The no-think baseline performs poorly, underscoring the necessity of deliberate reasoning in complex dialog states. DyBBT strikes a balance between these extremes by invoking costly reasoning only when cognitively justified, either due to under exploration or low confidence, leading to near optimal performance with moderate and targeted computational overhead. This leads to less efficient allocation of computational resources, as also reflected in human evaluation (Sec. 4.4).
E.9 Failure Mode Analysis and Limitations
While DyBBT demonstrates strong performance across benchmarks, we conducted a comprehensive failure mode analysis to understand its limitations in practical deployment scenarios. Through post-hoc analysis on 1000 dialogs of MultiWOZ with cross validation by three expert annotators, we quantitatively assessed the occurrence rates of different failure modes.
Category Description Rate Impact Level
Inaccurate Cognitive State Representation Handcrafted $\mathbf{c}_{t}$ fails to capture complex dialog dynamics like abrupt intent shifts 3.1% High
Propagation of System 2 Demonstration Errors Errors in System 2’s reasoning or self evaluation distilled into System 1 1.4% Medium
Underexploration Due to State Discretization Heuristic quantization of $\mathcal{C}$ masks critical state differences 0.7% Low
Total Failure Rate 5.2%
Table 11: Quantitative analysis of DyBBT failure modes on MultiWOZ dataset (N=1000 dialogs)
Table 11 presents the quantitative breakdown of failure modes, revealing that 94.8% of dialogs proceed without significant failures while only 0.3% exhibit multiple concurrent failure modes. The failure modes primarily occur in edge cases characterized by abrupt user intent shifts, complex cross domain dependencies, and non-standard user behaviors. These scenarios constitute inherently challenging “hard cases” that represent a minority in real world task oriented dialogs. The built-in safety mechanisms demonstrate substantial protective value: the Confidence Condition intercepts 76% of System 1’s low confidence errors, preventing catastrophic failures in uncertain states; Knowledge Distillation reduces System 2 invocation rate by 42% (Fig. 7), progressively mitigating error propagation risks; and human evaluation shows 88.7% alignment with expert judgment, far exceeding the random switching baseline. These builtin safety mechanisms demonstrate substantial protective value.
For the majority of commercial task oriented dialog scenarios, DyBBT’s current failure profile represents an acceptable risk given its significant performance advantages. However, in safety critical domains, the identified failure modes warrant additional safeguards. Our future work addresses these limitations through end-to-end learned cognitive representations, improved uncertainty calibration, and adaptive exploration mechanisms. These evolutionary improvements will further enhance DyBBT’s robustness while preserving its core architectural advantages for practical deployment.
E.10 Case Study
@@
E.10.1 Case 1: Successful Intervention due to High Epistemic Uncertainty
@@
  • Slot Dependency ($\rho_{t}$): 0.6 (Medium, ‘area’ and ‘name’ often co-occur in restaurant domain)
Meta-Controller Decision: The visitation count for this cognitive state region was low ($n_{t}(\mathbf{c}_{t})=12<\tau\sqrt{\log T}\approx 25$), triggering System 2 via the exploration condition. System 1’s confidence was medium ($p_{t}^{S1}=0.75>\kappa$). System 2 Intervention: System 2 performed multi-path reasoning. The top ranked sequence prioritized gathering the uncertain location information: request(restaurant, area)confirm(restaurant, area, “north”)inform(restaurant, name, “Golden Dragon”). Outcome: This strategy efficiently disambiguated the user’s intent. The dialog was successfully completed 6 turns later. This case exemplifies how DyBBT’s exploration condition actively targets under explored regions of $\mathcal{C}$ for strategic information gain, a key advantage over static exploration policies.
E.10.2 Case 2: Successful Intervention due to Low Aleatoric Confidence
@@
Meta-Controller Decision: The cognitive state was well explored ($n_{t}(\mathbf{c}_{t})=505$). However, System 1’s confidence was low ($p_{t}^{S1}=0.55<\kappa=0.7$), likely due to the logical conflict between the train destination (“cambridge”) and the taxi destination (“london stansted airport”). This triggered System 2 via the confidence condition. System 2 Intervention: System 2 identified the inconsistency. Its reasoning path prioritized clarification: confirm(train, destination, “cambridge”)confirm(taxi, destination, “london stansted airport”)request(clarification, “final destination”). Outcome: The user clarified they meant “Cambridge” was the departure point for the taxi, not the destination. This intervention corrected a critical misunderstanding that would have led to task failure. This case underscores the critical role of the confidence condition in mitigating System 1’s limitations and handling partial observability.
E.10.3 Case 3: Failure due to Cognitive State Misrepresentation
@@
Meta-Controller Decision: The state had medium visitation ($n_{t}(\mathbf{c}_{t})=162$) and System 1 was highly confident ($p_{t}^{S1}=0.92$) in its action to request(attraction, name). The meta-controller did not trigger System 2. Analysis of Failure: While the cognitive state suggested a routine information gathering context, it failed to capture the user had just finished a complex booking and was likely expecting a concise recommendation, not another request. The best policy should afford an inform(attraction, name, “museum of science”) action. Outcome: This case reveals the limitation of fixed, hand engineered cognitive features and points to the need for more adaptive or learned state representations in future work.
E.10.4 Summary and Limitations